[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Oct 13, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Oct 13 09:33:09 UTC 2016


HSU Shift Log for Thu Oct 13 2016 05:33:05 GMT-0400 (EDT)

Shift: A

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS/AWIPS2 issues:

Other Techincal issues: Check SDM logs for an issue with Gyre that could  
have been related to non-operational NHC requests on the supercomputer. I  
did not understand what could be done-- the SDM said he would make a log of  
the problem.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161013/04a8c73e/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list